Cr2O3+2Al=2Cr+Al2O3
Calculate the mass of aluminium needed to react with 1 tonne of chromium (iii) oxide.
1
Expert's answer
2012-10-01T10:05:37-0400
Cr2O3 + 2Al = 2Cr + Al2O3

M(Cr2O3) = 152 kg/kmol
M(Al) = 27
kg/kmol

n(Cr2O3) = 1000 kg / 152 kg/kmol = 6.579 kmol
n(Al) = 2n(Cr) =
6.579*2 = 13.158 kmol
m(Al) = n(Al)*M(Al) = 13.158 kmol * 27 kg/kmol = 355.3
kg